OpenStudy (anonymous):

what's the next number in the sequence 2 10 30 68

OpenStudy (anonymous):

68 + 24 = 92

OpenStudy (anonymous):

next term is 92 I think

OpenStudy (anonymous):

how is that possible

OpenStudy (anonymous):

plz give the reasoning

OpenStudy (anonymous):

hello there's no answer for my question

OpenStudy (anonymous):

is this a puzzle?

OpenStudy (anonymous):

no , its an I.O question

OpenStudy (anonymous):

I.O.?

OpenStudy (anonymous):

sorry, I.Q question

OpenStudy (anonymous):

ahh then there is something in the sequence that probably has nothing to do with adding or multiplying etc. some other connection like letters or words

OpenStudy (anonymous):

bet this has something to do with common difference. but we have so few numbers to work with

OpenStudy (anonymous):

10 - 2=8 30 - 10 = 20 68-30 = 38 now if you do it again you get 20-8=12 38-20=18 and one more time 18-12=6

OpenStudy (anonymous):

it would really be helpful to have one more term to see if the differences were 6, but who knows?

OpenStudy (anonymous):

i'm going to guess 130

OpenStudy (anonymous):

nothing striking me

OpenStudy (anonymous):

i am thinking along the lines of elecengineer, but instead of 92 add 92 to 68 to get 130. but i am just making things up. i thought of common differences but my method is stupid because i would need none more term.